1. (25 pts) Let F(x, y, z) = (2xy + 25)i + (4.r?y3 + 2yz?)j + (5.624 + 3y222)k and let C be the curve given parametrically byPlease explain all steps.

Thanks!
Answer #1

for ostsi solution F19,4.3) = (229* + 29) 1+ (98y?+ 2yz!)j + (5224+ 3y?z? Jé 04+) = (3t+1) ++?; + 5tk First we need the the v+ p (t) = 36+2+ ſ +5X Finally lets get the dot product taken care of (olp) .3142 18++678 + 937575 + 72+9+8+*+48 +*+ 500t